OpenStudy (anonymous):

What are the x- and y-intercepts of the quadratic equation y = x2 + 2x - 3?

OpenStudy (anonymous):

ie where x is zero and y is zero...

OpenStudy (anonymous):

\[(x+3)(x-1)=0\rightarrow x _{1}=-3 and x _{2}=1\]

OpenStudy (anonymous):

\[y=-3\]

OpenStudy (anonymous):

a. x-intercepts: (3, 0) and (-1, 0) y-intercept: (0, 3) b. x-intercepts: (-3, 0) and (1, 0) y-intercept: (0, -3) c. x-intercepts: (3, 0) and (-1, 0) y-intercept: (0, -3) d. x-intercepts: (-3, 0) and (1, 0) y-intercept: (0, 3)

OpenStudy (anonymous):

b. As @jeff1992 pointed out, x-intercepts are given by the roots of the polynomial. Those are x = -3 and x = 1. So, (-3,0) and (1,0) are the intercepts. The y intercept is when x = 0, therefore, y = 0^2 + 2*0 - 3 = -3. The point is (0,-3).